OpenStudy (anonymous):

if m∠T = 2x + 10 and m∠E = 3x+20, what is m∠H?

OpenStudy (anonymous):



OpenStudy (shamil98):

T + E = 180 T + H = 180 2x + 10 + 3x + 20 = 180 5x + 30 = 180 5x= 150 x= 30 so T + H = 180 2(30) + 10 + H = 180 60 + 10 + H = 180 H + 70 = 180 H = 110

OpenStudy (shamil98):

the angles are supplementary, which means they are equal to 180

OpenStudy (shamil98):

so you just solve for x and then solve for h .

OpenStudy (anonymous):

ahh. i see.. thank you :)